# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v1

@@ Abstract
LiDAR segmentation is crucial for autonomous driving systems. The recent range-view approaches are promising for real-time processing. However, they suffer inevitably from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose a simple yet powerful FRNet that restores the contextual information of the range image pixels with corresponding frustum LiDAR points. Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, which enables each point to extract more surrounding information via the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic prediction. Extensive experiments on four popular LiDAR segmentation benchmarks under various task setups demonstrate our superiority. FRNet achieves competitive performance while maintaining high efficiency. The code is publicly available.
1 Introduction
@@
Figure 1: Scalability comparisons among state-of-the-art LiDAR segmentation models on the SemanticKITTI [5] leaderboard. The size of the circular represents the number of model parameters. Our FRNet achieves competitive performance with current arts while maintaining satisfactory efficiency for real-time processing.
Real-time and accurate LiDAR semantic segmentation is crucial in autonomous driving, robotics, and other safety-critical applications [18]. Many recent approaches have made significant progress in this area, which represent LiDAR point clouds from different perspectives, including point view [48, 21], voxel view [47, 65], pseudo-image view [50, 35, 61], and multi-view representation [31, 54, 2], depending on different pre-processing techniques.
Point-view methods manipulate the origin point cloud to preserve raw spatial information, following the paradigm of PointNet++ [40]. Nevertheless, they require computationally expensive neighborhood searches to construct local neighborhoods, and the computation overhead is considerable for large scene point clouds. Voxel-view methods transform scatter points into regular voxel representations and apply 3D convolutions for label prediction, but the computation grows cubically when voxel resolution increases. Multi-view methods extract features from different representations separately and fuse these features from different levels for more accurate prediction. Consequently, the computation of all representations is also gathered. Albeit these representation methods achieve some satisfactory performance, their computation overhead hinders them from real-time online applications.
@@
Recently, pseudo-image methods propose a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [61, 1] and range-view [50, 51, 43, 4, 24]. Under such representations, some popular image segmentation approaches [33, 62] can be directly applied to such pseudo-images to balance the efficiency and accuracy of segmentation. However, the 3D-to-2D projection procedure tends to result in corrupted neighborhood contextual information, decreasing the precision. Detailed analysis of this problem can be summarized into two aspects.
Firstly, some 3D points are discarded in the projection procedure. Under such a paradigm, only the projected points are involved in convolutions, while the contextual information of discarded points is absent. Consequently, the pixel in the pseudo-image can hardly preserve the high-fidelity 3D geometric context. More importantly, these discarded 3D points still need to predict 3D labels but no features correspond to them. To reconstruct these labels, some post-processing techniques, such as CRF [22] and KNN [35], are used for these points based on 2D predictions. However, these techniques are unsupervised and fail to extract the local features of discarded points in a learnable manner, which significantly limits the improvement of pseudo-image methods.
Secondly, the pseudo-image suffers from many empty pixels due to the sparsity of LiDAR points. Under such a non-compact representation, the projected points will extract much noisy information as the convolutions go deeper. In summary, pseudo-image methods fail to preserve the contextual information of all the points, which is significant for LiDAR segmentation. So, it is important for points and pseudo-images to cooperate with each other. Although existing multi-view methods have successfully preserved the context of all points via fusing features of different representations, they are always computationally expensive.
@@
To this end, we study a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which can preserve the contextual information of all points through the cooperation of point and pseudo-image and still maintain high efficiency. Specifically, FRNet comprises three main components. To preserve all points and predict semantic labels for these points in an end-to-end manner, a Frustum Feature Encoder is used to group all points with the same frustum region into the corresponding range-view pixel in relation to the range image via several multi-layer perceptions. The point features can be pooled to represent the frustum region and formatted into a 2D representation, followed by traditional convolutions. Then, a Frustum-Point (FP) Fusion Module is introduced to update the features of each point hierarchically during each stage of the convolutions, which consists of frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. In this way, all the points extract larger local features based on the frustum region. Finally, a Fusion Head module is designed to utilize different level features to generate individual features for each point, so that the results can be predicted end-to-end without any post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvements among range-view methods while still maintaining high efficiency.
To alleviate empty pixels in 2D representation, we introduce a new range interpolation augmentation method that adds several potential points based on the information of surrounding points. Specifically, we first transform the point clouds into range images, as done in previous works. For empty pixels in the range images, we then calculate their potential coordinates when both the nearest left and right pixels have projected points as they often share similar attributes. With this augmentation, the 2D representation will have fewer empty pixels, allowing convolutions to learn more semantically coherent information.
The key contributions are summarized as follows:
  • We introduce a novel Frustum-Range representation for LiDAR segmentation. In this representation, each pixel in the 2D representation is treated as a corresponding frustum region in the point clouds. All points within the same frustum region are preserved for semantic prediction without the need for any post-processing techniques.
  • We propose a novel range interpolation augmentation that synthesizes possibly lost LiDAR points based on surrounding information, helping to fill empty pixels and learn more coherent range-view representations.
  • Extensive experiments across four LiDAR segmentation benchmarks demonstrate that FRNet can achieve the optimal trade-off between efficiency and accuracy.
2 Related Work
@@
Scalable LiDAR Segmentation. Although Point-view [21, 48], voxel-view [47, 65], and multi-view [31, 32] have achieved great success in LiDAR segmentation, they are often time-consuming and cannot achieve real-time processing. Recently, various methods proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation. SqueezeSeg [50] and SqueezeSegV2 [51] used the lightweight model SqueezeNet [22] to retain information. SqueezeSegV3 [53] introduced a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [35] integrated Darknet into SqueezeSeg [50] and proposed an efficient KNN post-processing for segmentation. RangeViT [4] and RangeFormer [24] introduced popular transformer blocks to extract local and global information of the range images. These methods only learn the features of projected points while ignoring the occluded points.
LiDAR Post-Processing. Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. Many of these methods follow the CRF [50] or KNN [35] approach to smooth the semantic results of occluded points. FIDNet [64] proposed a Nearest Label Assignment (NLA) method for label smoothing. Both of these techniques are unsupervised and can not extract the features of all points in a learnable manner. RangeFormer [24] created sub-clouds from full point clouds and inferred each subset in a supervised manner, but it still relies on 2D prediction results. In contrast, our method involves the frustum region and predicts the semantic label of each point directly without any information loss, making it a fully end-to-end approach.
@@
LiDAR Data Augmentation. Recent works have explored different data augmentation techniques on point clouds [7, 60], which can help improve performance when it comes to indoor scene understanding but has limited generalization to outdoor scenes. Mix3D [36] created new training samples by combining two scenes out of context. PolarMix [52] proposed mixing two LiDAR scenes based on yaw angles and cropping instances from one scene to another, while LaserMix [26] proposed exchanging labeled scans with unlabeled scans along the pitch axis for efficient semi-supervised learning. RangeFormer [24] presented a lightweight augmentation method that only operated on projected points in range images, disregarding occluded points. In this work, we prevent a novel augmentation technique based on our proposed frustum region, which preserves all points within the frustum region, encompassing techniques such as FrustumMix and Range-Interpolation.
@@
Figure 2: A pilot study on the performance degradation of post-processing in existing range-view methods [35, 12, 64, 8] on SemanticKITTI [5]. We choose various $K$ values as hyperparameters in KNN post-processing. Compared to their performance at 2D, a severe drop in performance occurs with different $K$ values.
3 FRNet: A Scalable LiDAR Segmentor
3.1 Pilot Study
@@
Figure 3: Architecture overview. The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features in the frustum region. 2) Frustum-Point (FP) Fusion Module is used to update per-point features hierarchically at each stage of the 2D backbone. 3) Fusion Head fuses different levels of features to predict final results.
@@
LiDAR segmentation aims to assign semantic labels to each point in the point clouds. A common approach is to perform operations directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [40]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this issue, many recent works propose projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Given point clouds with coordinates $(x,y,z)$ and the expected projected resolution $(H,W)$ of range images, each point can be mapped to its correspondence coordinates $(u,v)$ in the range image via the following transformation:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (1)
@@
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. However, such a projection process can cause the loss of occluded points and often requires post-processing techniques to reduce the blurry prediction of lost points [24].
@@
We conduct a pilot study on different range-view methods, including RangeNet++ [35], SalsaNeXt [12], FIDNet [64], and CENet [8], and compare their performance with different KNN post-processing parameters and 2D performance. As shown in Fig. 2, the post-processing procedures are often unsupervised, and the hyperparameters can significantly influence the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) that can directly predict semantic labels for each point while maintaining high efficiency.
3.2 Frustum-Range Architecture
@@
Given a point cloud $\mathbf{P}$ acquired by the LiDAR sensor, FRNet is expected to learn a feed-forward network $\Phi$ to predict semantic labels $\mathbf{Y}$ for each point:
$\mathbf{Y}=\Phi(\mathbf{P},\theta)~{},$ (2)
@@
where $\theta$ denotes the learnable parameters in the network. As shown in Fig. 3, FRNet comprises a Frustum Feature Encoder that carries out per-point feature extraction, a Frustum-Point Fusion Module that fuses per-point features and frustum features hierarchically, and a Fusion Head Module that predicts semantic labels for each point.
Frustum Feature Encoder. LiDAR segmentation involves predicting semantic labels for each point, so it is important to extract individual features for each point. Before converting point clouds into 2D representations, Frustum Feature Encoder first extracts per-point features via several MLPs, which is crucial for accurate label prediction. For points $\mathbf{P}_{i}=\{p_{i}^{1},p_{i}^{2},...,p_{i}^{N}\}$ in the $i$-th frustum region, the points are manually embedded with $(p_{i}^{k},p_{i}^{k}-p_{i}^{mean})$, where $p_{i}^{k}$ is the $k$-th point in the frustum, $N$ is the total number of points in the frustum region, and $p_{i}^{mean}=\frac{1}{N}\sum_{k=1}^{N}p_{i}^{k}$ represents the mean information of the frustum region. The per-point features are then obtained as follows:
$f_{i}^{k}=\text{MLP}([p_{i}^{k};p_{i}^{mean}])~{},$ (3)
@@
Frustum-Point Fusion Module. As shown in Fig. 4, the Frustum-Point (FP) Fusion Module comprises two components: 1) A Frustum-to-Point fusion that hierarchically updates per-point features, and 2) A Point-to-Frustum fusion to fuse individual features of each point within its corresponding frustum region. Specifically, for the point features $F_{p}$ and frustum features $F_{f}$, $F_{f}$ is back-project to the point-level features $F_{p}^{f}$ based on the projection index. This means that points within the same frustum region share the same frustum features. Then, the two features are concatenated together, followed by an MLP to update per-point features:
$\hat{F}_{p}=\text{MLP}([F_{p}^{f};F_{p}])~{}.$ (4)
Next, the updated per-point features $\hat{F}_{p}$ are pooled into the corresponding frustum region to further enhance the frustum features $F_{f}$ in a residual-attentive manner. The pooled features are denoted as $F_{f}^{p}$, which are calculated with the max operation based on the frustum region. The pooled features $F_{f}^{p}$ are first concatenated with $F_{f}$, followed by a simple convolutional layer to reduce the number of channels:
$F^{fuse}=g([F_{f}^{p};F_{f}],\theta_{1})~{},$ (5)
@@
where $g(\cdot)$ is the mapping function and $\theta_{1}$ is the learnable parameters. Finally, the fused features are used to further enhance the frustum features in a residual-attentive manner:
$\hat{F}_{f}=F_{f}+\sigma(h(F^{fuse},\theta_{2}))\odot F^{fuse}~{},$ (6)
@@
where $\sigma$ represents the sigmoid function, $h$ is the mapping function, $\theta_{2}$ is the learnable parameters, and $\odot$ indicates element-wise multiplication. The output features $\hat{F}_{f}$ are then used to extract larger local frustum features with subsequent convolutions.
Fusion Head Module. Since the aim of FRNet is to classify categories for each point, the most straightforward method is to directly use the updated per-point features after the FP Fusion Module for prediction. Although this can also achieve great performance, we observe that this is not the optimal solution. The output of our FP Fusion Module consists of two features: per-point features and frustum features represented in a 2D manner, and our prediction head expects to fully utilize both of these two features. Additionally, the per-point features generated from the Frustum Feature Encoder Module directly embed coordinate information, which can also provide more low-level geometric information. The Fusion Head Module progressively fuses three different level features. Firstly, the frustum features $F_{f}$ are mapped to point-level features $F_{f}^{p}$. It is worth noting that points in the same frustum regions share the same features at this stage. Then, we map the features $F_{f}^{p}$ with MLPs to ensure they share the same channels as point features. Finally, the add operation is applied to fuse the two features. The same operation is also used to fuse with low-level point features $F_{p}^{l}$. The overall operation can be formulated as:
$F_{out}=\text{MLP}(\text{MLP}(F_{f}^{p})+F_{p})+F_{p}^{l}~{}.$ (7)
Here, $F_{out}$ is used to generate the final semantic scores for each point.
Optimization. To improve the optimization of FRNet, we not only utilize the commonly used point-level supervised loss function but also supervise the frustum-level features with generated pseudo frustum labels. Typically, a frustum region may contain multiple points, and these points often do not belong to the same category. Based on this observation, we generate a pseudo label for each frustum in a statistical manner. We count the frequency of each category in the frustum region and choose the category with the highest frequency as the pseudo label for the frustum. Finally, the frustum-level features will be supervised using the generated pseudo labels. The overall loss can be calculated as:
$L=L_{p}+\lambda L_{f}~{},$ (8)
where $L_{p}$ and $L_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $L_{p}$ is calculated by cross-entropy loss and $L_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42].

### v2

@@ Abstract
LiDAR segmentation has become a crucial component in advanced autonomous driving systems. Recent range-view LiDAR segmentation approaches show promise for real-time processing. However, they inevitably suffer from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose FRNet, a simple yet powerful method aimed at restoring the contextual information of range image pixels using corresponding frustum LiDAR points. Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information via the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic prediction. Extensive experiments conducted on four popular LiDAR segmentation benchmarks under various task setups demonstrate the superiority of FRNet. Notably, FRNet achieves 73.3% and 82.5% mIoU scores on the testing sets of SemanticKITTI and nuScenes. While achieving competitive performance, FRNet operates 5 times faster than state-of-the-art approaches. Such high efficiency opens up new possibilities for more scalable LiDAR segmentation. The code has been made publicly available at https://github.com/Xiangxu-0103/FRNet.
Index Terms:
LiDAR Semantic Segmentation; Autonomous Driving; Real-Time Processing; Frustum-Range Representation
I Introduction
LiDAR segmentation, an crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4]. The principal challenge lies in devising a scalable LiDAR segmentation system that strikes a delicate balance between efficiency and accuracy, especially in resource-constrained operational scenarios [5, 6].
Existing LiDAR segmentation methods adopt various data representation perspectives, each with its own set of trade-offs in terms of accuracy and computational efficiency. The point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve raw spatial information, but entail computationally expensive neighborhood searches to construct local structures. This demand results in considerable computation overhead and limits their capability to deal with large-scale scene point clouds [13, 14, 10]. Sparse-voxel-based methods [15, 16, 17, 18] transform scattered LiDAR points into regular voxel representations and leverage popular sparse convolutions to extract features. However, this process also demands heavy computations, especially for applications that require high voxel resolutions [16, 15, 19]. Multi-view methods [20, 21, 22, 23] extract features from various representations to enhance prediction accuracy [24, 25, 26]. While this strategy can improve performance, it aggregates the computation required for all representations. Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [27, 23].
@@ I Introduction
Figure 1: A study on the scalability of state-of-the-art LiDAR segmentation models on the SemanticKITTI [28] leaderboard. The size of the circular representation corresponds to the number of model parameters. FRNet achieves competitive performance with current state-of-the-art models while still maintaining satisfactory efficiency for real-time processing.
@@
Recently, pseudo-image-based methods [29, 30, 31, 32, 33, 34, 35, 36] have appeared as a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [33, 34, 35] and range-view [29, 30, 31, 32], which are more computationally tractable. These representations enable the direct application of popular 2D segmentation approaches [37, 38, 39] to the pseudo-images, offering a promising solution for real-time LiDAR segmentation. However, the 3D-to-2D projection inevitably introduces corrupted neighborhood contextual information, which poses limitations on the further development of pseudo-image approaches [40]. The detailed analysis of this problem can be further summarized into two aspects as follows.
Firstly, a fundamental concern with pseudo-image methods is the inadvertent exclusion of certain 3D points during the conversion to 2D projections [41]. This leads to an incomplete representation where only the points that are projected are involved in subsequent convolutional operations. Such a scenario disregards the contextual information of the unprojected points, causing the pseudo-images to lose crucial high-fidelity 3D geometric context [42]. Furthermore, the omitted points, which still necessitate 3D label predictions, lack associated feature information, thereby necessitating post-processing to infer these labels. Unfortunately, current post-processing methods struggle to incorporate local features of these omitted points effectively, capping the potential advancements in pseudo-image approaches.
Secondly, the challenge with pseudo-images is exacerbated by the significant presence of empty pixels due to the inherent sparsity of LiDAR data. For instance, in the nuScenes dataset, nearly 60% of range image pixels are found to be void [43, 32]. This vacancy not only skews the scene representation but also compromises the quality of semantic segmentation. As the network processes deeper layers, these sparse pseudo-images can become riddled with noise, obstructing the extraction of meaningful patterns. In essence, although pseudo-image methods excel in terms of real-time processing capabilities, their reduced dimensional perspective can overlook crucial 3D details, leading to degraded scalability.
@@
Observing the above issues, in this work, we propose a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy. FRNet consists of three main components. Firstly, a frustum feature encoder (FFE) is utilized to group all points with the same frustum region into corresponding range-view pixels in relation to the range image using multiple multi-layer perceptrons (MLPs). This allows for the preservation of all points and the prediction of semantic labels for these points in an end-to-end manner. Subsequently, the point features are pooled to represent the frustum region and formatted into a 2D representation, which is then subjected to traditional convolutions. Secondly, a frustum-point (FP) fusion module is employed to efficiently update the hierarchical features of each point during each convolutional stage. This module includes frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. As a result, all points extract larger local features based on the frustum region. Finally, a fusion head (FH) module is designed to leverage features from different levels to generate individual features for each point, facilitating end-to-end prediction without the need for post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvement among range-view methods while still maintaining high efficiency.
Data augmentation plays an important role in training robust and generalizable models. Inspired by the success of mixing strategies [44, 45, 46], we introduce a novel mixing method dubbed FrustumMix, which operates based on frustum units. Specifically, we randomly divide point cloud scenes into several frustum regions along inclination and azimuth directions and swap the corresponding regions from different scenes, with the neighborhood of each frustum region coming from the other scene. FrustumMix provides a straightforward and lightweight method to adapt FRNet by focusing on the frustum region. Furthermore, to address the issue of empty pixels in the 2D representation, we propose RangeInterpolation to reconstruct the semantic surface based on surrounding point information. It first transforms the LiDAR point cloud into a range image, as done in previous works [29, 30, 31, 32]. For empty pixels in the range image, we then aggregate surrounding range information within a pre-defined window using an average pooling operation to generate a new point. RangeInterpolation is able to assist in creating a more compact 2D representation with fewer empty pixels, enabling convolutions to learn more semantically coherent information.
In summary, the key contributions of this work can be listed as follows:
  • We introduce a novel Frustum-Range representation for scalable LiDAR segmentation. In our framework, the point-level geometric information is integrated into the pseudo-image representation, which not only retains the complete geometric structure of the point cloud, but also takes advantage of the efficiency of the 2D network.
  • We propose two novel 3D data augmentation techniques to assist in training robust and generalizable models. FrustumMix generates more complex scenes by mixing two different scans, while RangeInterpolation reconstructs the semantic surfaces based on surrounding range information. These techniques help generate a compact 2D representation and learn more context-aware features.
  • Extensive experiments across four prevailing LiDAR segmentation benchmarks demonstrate our superiority. FRNet achieves $73.3\%$ and $82.5\%$ mIoU scores, respectively, on SemanticKITTI and nuScenes, while still maintaining promising scalability for real-time LiDAR segmentation.
II Related Work
II-A Scalable LiDAR Segmentation
@@
Although point-view [7, 8, 10, 11, 12], sparse-voxel-view [15, 16, 17, 18], and multi-view [20, 21, 22, 23, 47, 48] methods have achieved significant success in LiDAR segmentation, they often suffer from high computational costs and are unable to achieve real-time processing. Recently, various methods have been proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation [49]. SqueezeSeg [29] and SqueezeSegV2 [50] use the lightweight model SqueezeNet [51] to retain information. SqueezeSegV3 [52] introduces a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [30] integrates DarkNet into SqueezeSeg [29] and proposes an efficient KNN post-processing for segmentation. RangeViT [31] and RangeFormer [32] introduce transformer blocks to extract both local and global information from the range image. However, these methods only involve projected points and fail to extract contextual information for discarded points. To address this problem, WaffleIron [11] proposes to mix point and image features with residual connections. Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predicting over the entire point cloud. In this work, we propose to integrate point and image features efficiently, preserving both 3D geometric information from raw point features and efficiency from 2D convolutions, without the need for any pre- or post-processing techniques, which yields superior scalability for real-time LiDAR segmentation.
II-B LiDAR Data Augmentation
@@
Recent works have explored various data augmentation techniques on point clouds [53, 54], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes. Mix3D [46] creates new training samples by combining two scenes out of context, which requires a larger memory cost. PolarMix [45] proposes to mix two LiDAR scenes based on azimuth angles and cropping instances from one scene to another, while LaserMix [44] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning. RangeFormer [32] presents a lightweight augmentation method that operates only on projected points in range images, disregarding occluded points. In this work, we introduce augmentation techniques aimed at training a robust and scalable model for LiDAR segmentation, including FrustumMix and RangeInterpolation.
II-C LiDAR Post-Processing
Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. SqueezeSeg [29] employs the traditional conditional random field (CRF) as a recurrent neural network (RNN). RangeNet++ [30] introduces a fast K-Nearest-Neighbor (KNN) search to aggregate final results across the entire point cloud. FIDNet [55] uses a Nearest Label Assignment (NLA) method for label smoothing. RangeFormer [32] divides full point clouds into sub-clouds and infers each subset in a supervised manner. However, these methods heavily rely on 2D results and fail to extract point features over entire points in a learnable manner. In contrast, KPRNet [56] and RangeViT [31] incorporate a KPConv [7] block to extract point features in an end-to-end manner. Nevertheless, this approach requires substantial computational resources to search neighborhoods in 3D space, which hinders them from real-time processing. In this work, we partition points into frustum regions and extract local features of points in a 2D manner, avoiding time-consuming neighbor searches. Based on the learned point features, FRNet predicts results over the entire point cloud in an end-to-end manner with high efficiency.
III FRNet: A Scalable LiDAR Segmentor
@@
Figure 2: Pilot study on the performance degradation of post-processing in existing range-view methods [30, 57, 55, 58] on the val set of SemanticKITTI [28]. We choose various $K$ values as hyperparameters in KNN post-processing. Compared to their performance at 2D (i.e., the range image), a severe drop in performance occurs with different $K$ values.
In this section, we first conduct a pilot study on the most popular range-view LiDAR segmentation methods, wherein we observe a significant impact of KNN post-processing on performance (Sec. III-A). We then elaborate on the technical details of frustum-range architecture (Sec. III-B) and frustum-range operators (Sec. III-C). The overall pipeline of the proposed FRNet framework is depicted in Fig. 3.
III-A Pilot Study
@@
Figure 3: Architecture overview. The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features within the frustum region. 2) Frustum-Point (FP) Fusion Module updates per-point features hierarchically at each stage of the 2D backbone. 3) Fusion Head fuses different levels of features to predict final results.
@@
LiDAR segmentation aims to assign semantic labels to each point in the point cloud [59]. A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [60]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this problem, many recent works have proposed projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})f^{-1}H]}~{},$ (1)
@@
where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, and $\phi_{\text{up}}$ are $\phi_{\text{down}}$ are the inclination angles at the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image. However, such projection often results in the loss of occluded points and requires post-processing techniques to reconstruct the entire semantic information [32].
@@
We conducted a pilot study on various range-view methods, including RangeNet++ [30], SalsaNext [57], FIDNet [55], and CENet [58]. We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with various nearest neighbor parameters, as well as their performance on the 2D range images. As illustrated in Fig. 2, the post-processing procedures are often unsupervised, and the choice of hyperparameters can significantly impact the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) capable of directly predicting semantic labels for each point while maintaining high efficiency.
III-B Frustum-Range Architecture
@@
Problem Definition. Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\hat{Y}}$ for each individual point as follows:
$\mathcal{\hat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
@@
where $\theta$ represents the learnable parameters within the network. As depicted in Fig. 3, FRNet consists of three key components: 1) a frustum feature encoder responsible for per-point feature extraction; 2) a frustum-point fusion module that efficiently fuses hierarchical point features and frustum features; 3) a fusion head module that integrates point features at different levels for accurate semantic prediction.
Operator Representation. Let $\text{Flat}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{H\times W\times C}$ denotes the function that projects point features onto the frustum image plane with resolution $(H,W)$. Within each frustum region, we apply a max-pooling function to obtain the frustum features. Let $\text{MLP}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{N\times D}$ represents the MLPs that take $C$-dimensional point features as input and output $D$-dimensional point features. Let $\text{Inflat}(\cdot):\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{N\times C}$ be the operation that back-projects frustum features to point features. Notably, points falling into the same frustum region share the same features under this operation.
Frustum Feature Encoder. In LiDAR segmentation, the accurate prediction of semantic labels for each point requires the extraction of individual features for each point. Preceding the conversion of the point cloud into a 2D representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs. This process is crucial for accurate label prediction. Firstly, We group the point cloud to yield a set of $M$ frustum regions $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$ according to Eq. 1, where $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$, consisting of $N_{i}$ points. To explicitly embed the frustum structure, we leverage an average pooling function to obtain the cluster points $\widetilde{\mathcal{P}}=\{\widetilde{\mathcal{P}}_{1},\widetilde{\mathcal{P}}_{2},...,\widetilde{\mathcal{P}}_{M}\}$, where $\widetilde{\mathcal{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$. Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\widetilde{\mathcal{P}}$ within each frustum region. The per-point features are obtained as follows:
$\mathcal{F}_{p}^{0}=\text{MLP}([\mathcal{P};\mathcal{P}-\widetilde{\mathcal{P}}])~{},$ (3)
@@
Frustum-Point Fusion Module. Following prior works [55, 58, 32], we utilize a 2D backbone composed of multiple convolution blocks that efficiently extract hierarchical frustum features. Each stage block is accompanied by a frustum-point fusion module, facilitating the incremental update of point and frustum features. As illustrated in Fig. 4, the frustum-point fusion module consists of two essential components: 1) A frustum-to-point fusion that enables the hierarchical update of per-point features, and 2) A point-to-frustum fusion that fuses individual features of each point within its corresponding frustum region. The $i$-th frustum-point fusion module takes $\mathcal{F}_{p}^{i-1}$ and ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ as inputs, where ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is the output of the 2D convolution block that takes $\mathcal{F}_{f}^{i-1}$ as input. To efficiently extract context information of the point, we employ a frustum-to-point fusion block to integrate local features of the frustum region into points. Firstly, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index. Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update per-point features, avoiding the high-computation in neighborhood searches:
$\mathcal{F}_{p}^{i}=\text{MLP}([\text{Inflat}(\widetilde{\mathcal{F}}_{f}^{i-1});\mathcal{F}_{p}^{i-1}])~{}.$ (4)
As discussed above, the 2D representation is non-compact due to the sparsity of the point cloud. The convolution block inevitably introduces noisy information for the empty frustum, leading to compact feature representation. To preserve the sparsity attributes of the 2D representation, we design a point-to-frustum fusion module to fuse the sparse and dense frustum features. Specifically, the updated point features $\mathcal{F}_{p}^{i}$ are first pooled into the corresponding frustum region to yield a non-compact 2D representation. Subsequently, the non-compact frustum feature is concatenated with the compact frustum feature ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ and passed through a simple convolutional layer to reduce the number of channels:
$\mathcal{F}_{\text{fuse}}^{i}=\text{Conv}([\text{Flat}(\mathcal{F}_{p}^{i});\widetilde{\mathcal{F}}_{f}^{i-1}])~{},$ (5)
@@
where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and activation function. Finally, the fused features are utilized to further enhance the frustum features in a residual-attentive manner:
$\mathcal{F}_{f}^{i}=\widetilde{\mathcal{F}}_{f}^{i-1}+\sigma(h(\mathcal{F}_{\text{fuse}}^{i},\theta))\odot\mathcal{F}_{\text{fuse}}^{i}~{},$ (6)
@@
where $\sigma$ represents the sigmoid function, $h(\cdot)$ is a linear function with trainable parameters $\theta$, and $\odot$ indicates element-wise multiplication.